Find two consecutive integers such that 3 times the first integer plus 2 times the second integer is equal to 107.
Answer by checkley77(12844)

LET X= THE FIRST INTEGER
LET X+1=THE NEXT NUMBER
3X+2(X+1)=107
3X+2X+2=107
5X=107-2
X=105
X=105/5
X=21 IS THE FIRST INTEGER.
21+1=22 IS THE NEXT NUMBER.
PROOF
3*21+2*22=107
63+44=107
107=107